OpenStudy (anonymous):

for the given function, find the points on the graph at which the tangent line has slope 1. y=\frac{ 1 }{ 3 }x^3-2x^2+4x+1

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 3 }x^3-2x^2+4x+1\]

OpenStudy (anonymous):

Find points where \[\frac{dy}{dx}=1\]

OpenStudy (anonymous):

ok so would it be this? \[1*\frac{ d }{ dx }=\frac{ 1 }{ 3 }x^3*\frac{ d }{ dx }-2x^2*\frac{ d }{ dx }+4x*\frac{ d }{ dx }+1\]

OpenStudy (anonymous):

Why do you have d/dx multiplied into the equation like that? I mean to find the derivative of y and set it equal to 0.

OpenStudy (anonymous):

i dont know why i did that.